# Emergent Berezinskii-Kosterlitz-Thouless deconfinement in super-Coulombic plasmas

### v1

@@ Acknowledgements.
Acknowledgments – We thank Pradyumna Belgaonkar for useful discussions. S.G. acknowledges support from the Israel Science Foundation (ISF) Grant no. 586/22 and the US–Israel Binational Science Foundation (BSF) Grant no. 2020264. Computational resources were provided by the Intel Labs Academic Compute Environment and the Fritz Haber Center for Molecular Dynamics, The Hebrew University of Jerusalem. L.R. thanks John Toner for discussions, feedback on the manuscript, and collaboration on Ref. 18. L.R. acknowledges financial support by the Simons Investigator Award from the James Simons Foundation, and thanks Physics Department of The Hebrew University of Jerusalem for hospitality during a visit when this project was conceived.
@@ I Berezinskii-Kosterlitz-Thouless transition in a Coulomb Gas
As a validation step for our test charge-based approach, we study the BKT transition in a bare Coulomb gas. We compare this result with the standard method of calculating the effective Coulomb coupling using the charge density-density correlations given by the following relation arrived at using linear response theory [25],
@@
In Fig. S1, we compare the effective Coulomb coupling, extracted from both the density-density correlations (Fig. S1a) and our test charge based approach (Fig. S1b). Indeed, we find that in both cases, the coupling flows to zero in the thermodynamic limit beyond a critical temperature $T\gtrsim T_{\text{BKT}}$, marking a deconfinement transition. We estimate $T_{\text{BKT}}\approx 0.19$ in both cases, using the crossing with the BKT line $\kappa=4T$. However, we note that the precise finite system size results differ between the two methods. This can be attributed to the need to take the $k\to 0$ limit in Eq. S1, which is only justified at infinite system sizes. Our test charge approach, on the other hand, shows faster convergence to the thermodynamic result.
@@ II Signature of deconfinement in super-Coulombic density-density correlations
Figure S2: The charge density-density correlation divided by $Tk^{2}$ in the long wavelength limit as a function of temperature for a $\sigma=0.25$, $\rho=0.01$ super-Coulombic gas. The dashed line marks the thermodynamic BKT temperature extracted from Fig. 3 in the main text. To alleviate boundary effects due to the finite Monte Carlo box size, the low momentum limit is approached by taking $k=3\times 2\pi/L$.
@@
In Fig. S2 we plot $\lim_{k\to 0}\langle n(k)n(-k)/(k^{2}T)\rangle$ for a $\sigma=0.25$, $\rho=0.01$ super-Coulombic gas as a function of temperature for different system sizes. In the bound phase, for $T\leq T_{\text{BKT}}$, we see our observable takes a constant value for all system sizes. Since when approaching the limit $k\to 0$ with a fixed particle density $\rho$, the lowest discrete momentum $k\propto L^{-1}$ has an implicit scaling with system size, the $N$ independence of our observable for $T<T_{\text{BKT}}$ implies that $\langle n(k)n(-k)\rangle\sim k^{2}$ in the bound phase. By contrast, the peeling off of the different curves for $T\gtrsim T_{\text{BKT}}$ suggests an abrupt change in the functional form of $\langle n(k)n(-k)\rangle$ across $T_{\text{BKT}}$, signaling charge deconfinement.
@@ IV The Crossover Lengthscale
Figure S5: (a) The effective test potential in a $\sigma=0.25$ gas with a very low density $\rho=0.0002$ at three different temperatures. The vertical dashed lines show the points where the effective potentials deviate from the bare potential by more than one percent. (b) The thermodynamic coupling constant (left axis) $\kappa_{\infty}$ as extracted in Fig. 4a of the main text and the associated crossover scale upto an $O(1)$ calculated using the relation $\xi_{c}\times O(1)=\kappa^{1/\sigma}_{\infty}$ (right axis).
In this section, we study the super-Coulombic to Coulomb crossover length scale. Due to the observed long-tailed nature of the crossover, the true Coulomb phase is only achieved when $\kappa_{L}$ saturates. Here, we study the length scale where screening effects become significant. To make a direct measurement of $\xi_{c}$, we consider a super-Coulombic gas at extremely low density where there is a clearer separation of scales between $\rho^{-1/2}\gg\xi_{c}\gg a$ in Fig. S5a. We define $\xi_{c}$ as the distance where the effective potential deviates from it bare form by more than one percent, i.e., when $|V_{\text{eff}}(r)/n^{2}_{t}-r^{\sigma}|/r^{\sigma}\geq 0.01$. In Fig. S5a we use this metric to visually showcase $\xi_{c}$ at three different temperatures. It is evident that $\xi_{c}$ increases at lower temperatures.
Alternatively, using the relation $\xi_{c}\times O(1)=(\kappa_{\infty})^{1/\sigma}$ from [18], we extract $\xi_{c}$ upto an $O(1)$ constant in Fig. S5b from the extrapolated $\kappa_{\infty}$ values in Fig. 4a of the main text. We see that $\xi_{c}$ shows a divergence similar to $\kappa_{\infty}$ at lower temperatures.
@@ V Simulation Details
To benchmark our simulation, we compare our MCMC results against a “random dart” quasi-Monte Carlo calculation. Here, a configuration $c$ is drawn using quasi-random Sobol sequences [27] and assigned a weight $e^{-\beta H(c)}$. The expectation values of observables are then computed as $\langle O\rangle=\frac{\sum_{c}O(c)e^{-\beta H(c)}}{\sum_{c}e^{-\beta H(c)}}$. A benchmark comparison between our MCMC and the random dart quasi-Monte Carlo methods is shown in Fig. S6.
@@ VI Extracting the effective potential
In this section, we detail our numerical approach behind computing the effective test charge potential introduced in Eq. 2 of the main text. To that end, it is illuminating to rewrite Eq. 2 as:
$e^{-\beta V_{\text{eff}}(x_{1},x_{2})}=\langle e^{-\beta(H_{C}+H_{I})}\rangle_{H_{M}}=\frac{\int_{\{y_{i}\}}e^{-\beta H}}{\int_{\{y_{i}\}}e^{-\beta H_{M}}}=\frac{\mathcal{Z}(x_{1},x_{2})}{\mathcal{Z}_{0}},$ (S2)
where, $\mathcal{Z}(x_{1},x_{2})$ and $\mathcal{Z}_{0}$ are partition functions of the system in the presence and absence of test charges respectively.
One way to compute the effective potential in Eq. S2 is to consider fictitious test charges, i.e., test charges whose interaction energy doesn’t feature in the detailed balance calculations of the simulation. In such a case, the observable $\langle e^{-\beta(H_{C}+H_{I})}\rangle$ gets averaged over the distribution $e^{-\beta H_{M}}/\mathcal{Z}_{0}{}$. Note that this allows us to compute the entire potential $V_{\textrm{eff}}(r)\forall r$ in a single measurement pass since a given medium realization is not tied to any particular configuration of test charges.
@@
$G=\sum_{\{r\}}\mathcal{Z}_{r},$ (S3)
where, $\{\mathcal{Z}_{r}\}$ is the partition function of a configuration with test charges fixed at two locations in the bulk separated by a distance $r$. $\mathcal{Z}_{0}$ is again the partition function of the medium in the absence of any test charge. Using Eq. S2 and Eq. S3, the effective potential at a distance $r$ can be computed from the sampled histogram of counts registered for the different $r$ sectors of G as:
$e^{-\beta V_{\text{eff}}(r)}=\frac{\mathcal{Z}_{r}}{\mathcal{Z}_{0}}=\left\langle\frac{\delta_{r,r^{\prime}}}{\delta_{0,r^{\prime}}}\right\rangle_{G}.$ (S4)
This sampling approach is similar to the one employed in a different problem in [28]. Since both our methods are equivalent, their comparison presented in Fig. S7 serves as a benchmark. Due to its better scalability, the fictitious charge approach is employed for the data in the main text.
@@ VIII Effect of hard-sphere defects
When studying the form of the potential at short distances, like in Fig. 2 and Fig. S5a, we simulate the medium with defects. For other cases, when computing the long-wavelength Coulomb coupling or charge density-density correlations, both choices of mediums yield identical results.

### v2

@@ Acknowledgements.
Acknowledgments – We thank Pradyumna Belgaonkar for useful discussions. S.G. acknowledges support from the Israel Science Foundation (ISF) Grant no. 586/22 and the US–Israel Binational Science Foundation (BSF) Grant no. 2020264. Computational resources were provided by the Intel Labs Academic Compute Environment and the Fritz Haber Center for Molecular Dynamics, The Hebrew University of Jerusalem. L.R. thanks John Toner for discussions, feedback on the manuscript, and collaboration on Ref. Radzihovsky_2024. L.R. acknowledges financial support by the Simons Investigator Award from the James Simons Foundation, and thanks Physics Department of The Hebrew University of Jerusalem for hospitality during a visit when this project was conceived.
Data Availability – The data that supports our results can be found in Ref. data.
@@ I Berezinskii-Kosterlitz-Thouless Transition in a Coulomb Gas
As a validation step for our test charge-based approach, we study the BKT transition in a bare Coulomb gas. We compare this result with the standard method of calculating the effective Coulomb coupling using the charge density-density correlations given by the following relation arrived at using linear response theory [Orkoulas_1996],
@@
In Fig. S1, we compare the effective Coulomb coupling, extracted from both the density-density correlations (Fig. S1a) and our test charge-based approach (Fig. S1b). Indeed, we find that in both cases, the coupling flows to zero in the thermodynamic limit beyond a critical temperature $T\gtrsim T_{\text{BKT}}$, marking a deconfinement transition. We estimate $T_{\text{BKT}}\approx 0.19$ in both cases, using the crossing with the BKT line $\kappa=4T$. However, we note that the precise finite system size results differ between the two methods. This can be attributed to the need to take the $k\to 0$ limit in Eq. S1, which is only justified at infinite system sizes. Our test charge approach, on the other hand, shows faster convergence to the thermodynamic result.
@@ II Signature of deconfinement in Super-Coulombic density-density correlations
Figure S2: The charge density-density correlation divided by $Tk^{2}$ in the long wavelength limit as a function of temperature for a $\sigma=0.25$, $\rho=0.01$ super-Coulombic gas. The dashed line marks the thermodynamic BKT temperature extracted from Fig. 3a. To alleviate boundary effects due to the finite Monte Carlo box size, the low momentum limit is approached by taking $k=3\times 2\pi/L$.
@@
In Fig. S2, we plot $\lim_{k\to 0}\langle n(k)n(-k)/(k^{2}T)\rangle$ for a super-Coulombic gas as a function of temperature for different system sizes. In the bound phase, for $T\leq T_{\text{BKT}}$, we see our observable takes a constant value for all system sizes. Since when approaching the limit $k\to 0$ with a fixed particle density $\rho$, the lowest discrete momentum $k\propto L^{-1}$ has an implicit scaling with system size, the $N$ independence of our observable for $T<T_{\text{BKT}}$ implies that $\langle n(k)n(-k)\rangle\sim k^{2}$ in the bound phase. By contrast, the peeling off of the different curves for $T\gtrsim T_{\text{BKT}}$ suggests an abrupt change in the functional form of $\langle n(k)n(-k)\rangle$ across $T_{\text{BKT}}$, signaling charge deconfinement.
@@ IV The Crossover Lengthscale
Figure S5: (a) The effective test potential in a $\sigma=0.25$ gas with a very low density $\rho=0.0002$ at three different temperatures. The vertical dashed lines show the points where the effective potentials deviate from the bare potential by more than one percent. (b) The thermodynamic coupling constant (left axis) $\kappa_{\infty}$ as extracted in Fig. 4a and the associated crossover scale upto an $O(1)$ constant calculated using the relation $\xi_{c}\times O(1)=\kappa^{1/\sigma}_{\infty}$ (right axis).
We now study the super-Coulombic to Coulomb crossover length scale. Due to the observed long-tailed nature of the crossover, the true Coulomb phase is only achieved when $\kappa_{L}$ saturates. Here, we focus on the length scale where screening effects become significant. To make a direct measurement of $\xi_{c}$, we consider a super-Coulombic gas at extremely low density where there is a clearer separation of scales between $\rho^{-1/2}\gg\xi_{c}\gg a$ in Fig. S5a. We define $\xi_{c}$ as the distance where the effective potential deviates from it bare form by more than one percent, i.e., when $|V_{\text{eff}}(r)/n^{2}_{t}-r^{\sigma}|/r^{\sigma}\geq 0.01$. In Fig. S5a, we use this metric to visually showcase $\xi_{c}$ at three different temperatures. It is evident that $\xi_{c}$ increases at lower temperatures.
Alternatively, using the relation $\xi_{c}\times O(1)=(\kappa_{\infty})^{1/\sigma}$ from Ref. Radzihovsky_2024, we extract $\xi_{c}$ upto an $O(1)$ constant in Fig. S5b from the extrapolated $\kappa_{\infty}$ values in Fig. 4a. We see that $\xi_{c}$ shows a divergence similar to $\kappa_{\infty}$ at lower temperatures.
@@ V Simulation Details
To benchmark our simulation, we compare our MCMC results against a “random dart” quasi-Monte Carlo calculation. Here, a configuration $c$ is drawn using quasi-random Sobol sequences [Sobol_1967] and assigned a weight $e^{-\beta H(c)}$. The expectation values of observables are then computed as $\langle O\rangle=\frac{\sum_{c}O(c)e^{-\beta H(c)}}{\sum_{c}e^{-\beta H(c)}}$. A benchmark comparison between our MCMC and the random dart quasi-Monte Carlo methods is shown in Fig. S6.
@@ VI Extracting the effective potential
We now detail our numerical approach to computing $V_{\text{eff}}$ introduced in the main text. To that end, it is illuminating to rewrite Eq. 2 as:
$e^{-\beta V_{\text{eff}}(x_{1},x_{2})}$ $=\langle e^{-\beta(H_{C}+H_{I})}\rangle_{H_{M}}$ (S2)
$=\frac{\int_{\{y_{i}\}}e^{-\beta H}}{\int_{\{y_{i}\}}e^{-\beta H_{M}}}=\frac{\mathcal{Z}(x_{1},x_{2})}{\mathcal{Z}_{0}},$
where, $\mathcal{Z}(x_{1},x_{2})$ and $\mathcal{Z}_{0}$ are partition functions of the system in the presence and absence of test charges, respectively.
One way to compute the effective potential in Eq. S2 is to consider fictitious test charges, i.e., test charges whose interaction energy doesn’t feature in the detailed balance calculations of the simulation. In such a case, the observable $\langle e^{-\beta(H_{C}+H_{I})}\rangle$ gets averaged over the distribution $e^{-\beta H_{M}}/\mathcal{Z}_{0}{}$. Note that this allows us to compute the entire potential $V_{\text{eff}}(r)\forall r$ in a single measurement pass since a given medium realization is not tied to any particular configuration of test charges.
@@
$\mathcal{G}=\sum_{\{r\}}\mathcal{Z}_{r},$ (S3)
where, $\mathcal{Z}_{r}$ is the partition function of a configuration with test charges fixed at two locations in the bulk separated by a distance $r$. $\mathcal{Z}_{0}$ is again the partition function of the medium in the absence of any test charge. Using Eq. S2 and Eq. S3, $V_{\text{eff}}(r)$ can be computed from the sampled histogram of counts registered for the different $r$ sectors of the distribution $P(r)=\mathcal{Z}_{r}/\mathcal{G}$ as:
$e^{-\beta V_{\text{eff}}(r)}=\frac{\mathcal{Z}_{r}}{\mathcal{Z}_{0}}=\frac{\langle\delta_{r,r^{\prime}}\rangle_{r^{\prime}}}{\langle\delta_{0,r^{\prime}}\rangle_{r^{\prime}}}.$ (S4)
This sampling approach is similar to the one employed for a different problem in Ref. De_2024. Since both our methods are equivalent, their comparison presented in Fig. S7 serves as a benchmark. Due to its better scalability, the fictitious charge approach is employed for the data in the main text.
@@ VIII Effect of hard-sphere defects
When studying the form of the potential at short distances, like in Fig. 2, we simulate the medium with defects. For other cases, when computing the long-wavelength Coulomb coupling or charge density-density correlations, both choices of mediums yield identical results.
IX Boundary Effects
Figure S10: (a) The form of the potential at $\rho=0.01$ and $T=0.075$ up to a fixed distance in the bulk for different system sizes, contrasted against the bare potential. (b) The ratio of potentials between successive system sizes.
For systems with long-range interactions, as studied in this work, boundary effects can significantly affect emergent interactions in the bulk. To subvert this, as described in the main text, we choose an $L/3$ sized boundary buffer around the bulk region of study that approaches infinity in the limit of infinite box size. To show that this approach indeed systematically eliminates boundary effects, we study the potential up to a fixed distance $r_{\max}$ but for increasing system sizes in Fig. S10. Since the particle density $\rho$ is constant, increasing $N$ increases the box size $L$, pushing the region of study deeper into the bulk. In Fig. S10a we see that the potentials show minimal finite size effects, with successive system sizes converging to the same potential as $N$ approaches infinity. In contrast, the bare potential shows a drastically different functional form. In Fig. S10b we plot the ratio between the potentials obtained from systems of size $N$ and $N/2$. We see that for all $r$, this ratio systematically goes to $1$ as $N$ approaches the thermodynamic limit.